OpenStudy (anonymous):

If f(x) = -4x - 1, find f(-2)

OpenStudy (raden):

it means : f(-2) = -4(-2) - 1 = 8 - 1 = ....

Directrix (directrix):

Look at the function -->f(x) = -4 x - 1. What does f do to x? It multiplies x by negative four and then subtracts one.

Directrix (directrix):

For x -2, the function does the following: It multiplies (-2) by negative four and then subtracts one. So, what is the output?:

OpenStudy (anonymous):

so it would be negative 7

Directrix (directrix):

>so it would be negative 7 No. Recalculate. f(-2) = -4 * (-2) - 1 f(-2) = +8 - 1 f(-2) = ?

Directrix (directrix):

What is 8 minus 1?